# GradNetOT: Learning Optimal Transport Maps with GradNets

### v1

@@ Abstract
Monotone gradient functions play a central role in solving the Monge formulation of the optimal transport problem, which arises in modern applications ranging from fluid dynamics to robot swarm control. When the transport cost is the squared Euclidean distance, Brenier’s theorem guarantees that the unique optimal map is the gradient of a convex function, namely a monotone gradient map, and it satisfies a Monge-Ampère equation. In [1, 2], we proposed Monotone Gradient Networks (mGradNets), neural networks that directly parameterize the space of monotone gradient maps. In this work, we leverage mGradNets to directly learn the optimal transport mapping by minimizing a training loss function defined using the Monge-Ampère equation. We empirically show that structural bias of mGradNets facilitates the learning of optimal transport maps and employ our method for a robot swarm control problem.
Index Terms:
Optimal Transport, Gradient Networks, Monge-Ampère, Brenier’s Theorem, Physics Informed Neural Networks (PINNs)
@@ I Introduction
The optimal transport problem originated with the work of Gaspard Monge in 1781, who was interested in a fundamental problem that arises when constructing fortifications: how can one transfer molecules of soil from piles to holes while minimizing the cost of transfer [3]? Since then, Monge’s optimal transport problem and Leonid Kantorovich’s relaxation [4, 5] have been phrased in terms of probability densities and finding an optimal transport map: how can a source probability distribution be mapped to a target probability distribution in a way that minimizes the transport cost?
These formulations of optimal transport, along with several newer variations, have found applications spanning not only statistics and logistics, but also signal and image processing, robotics, machine learning, fluid dynamics, economics, geology, astronomy, and beyond. Economic applications of optimal transport include works showing the equivalence between optimal transport, stable matching, and modeling hedonic price equilibria, which capture relationships between prices and product characteristics [6]. Developments concerning optimal transport on Riemannian manifolds have been used to model continental drift by describing ground mass distributions and geodesic transport paths over time [7]. Recently, works on Monge-Ampère gravity employ optimal transport maps to model the distribution of mass throughout the universe and the formation of cosmological structures [8].
In our simulations, we illustrate an example of optimal transport used to control a robot swarm. As developments in reliable unmanned aerial vehicles (UAVs) lead to widespread availability and use, controlling large swarms of robots or drones is becoming a significant problem [9]. Swarms have applications ranging from critical natural disaster rescue missions to intricate and artistic light shows at festivals. Movements of large swarms can be modeled as time-evolving spatial probability distributions where an optimal transport map can guide the swarm from an initial distribution to a desired target distribution while minimizing travel cost [10].
@@
Exploiting Brenier’s link between optimal transport the Monge-Ampère equation is nontrivial because the Monge-Ampère equation is highly non-linear and generally difficult to solve using standard techniques like calculus of variations [12]. While some works, such as [8], rely on additional assumptions like spherical symmetry to recover an analytic solution, the majority of the literature uses numerical approaches like discrete-discrete entropic regularization and semi-discrete optimal transport [13]. In this paper, we focus on the intersection of Brenier’s result and neural networks to solve Monge optimal transport problems.
We propose GradNetOT: a method that solves Monge optimal transport problems by leveraging monotone gradient networks (mGradNets) [2]. Using an mGradNet to parameterize the optimal transport map, we search through the space of all gradients of convex potential functions. To guide our search, we explicitly use the Monge-Ampère equation in our optimization objective.
@@ II Background
[…]
@@
$f_{\#}\mu(A)=\mu(f^{-1}(A))=\mu(\{x\in\mathcal{X}:f(x)\in A\})$ (1)
where $f^{-1}(A)$ corresponds to the preimage of $A$ under $f$. Intuitively, (1) computes the probability of set $A$ in the target space $\mathcal{Y}$ by measuring the probability of the set in the source space $\mathcal{X}$ whose image under $f$ is $A$. Additionally, $f$ is called a diffeomorphism if it is a bijection and both $f$ and $f^{-1}$ are differentiable.
@@ III Problem Statement
Optimal transport involves finding the most efficient way to transform one probability distribution into another. The distributions can be thought of as describing how mass (representing probability) is spread over a space. Efficiency here is measured by a predefined cost function.
Formally, let $\mathcal{X}$ and $\mathcal{Y}$ be spaces equipped with respective probability measures $\mu$ and $\nu$. Let $c:\mathcal{X}\times\mathcal{Y}\to[0,\infty)$ be a cost function, where $c(x,y)$ denotes the cost of moving a unit of mass (probability) from $x\in\mathcal{X}$ to $y\in\mathcal{Y}$. The optimal transport problem is then to find a map $T:\mathcal{X}\to\mathcal{Y}$ that pushes forward $\mu$ to $\nu$ (i.e., $T_{\#}\mu=\nu$), and minimizes the total transport cost:
[…]
The formulation in (2) is known as the Monge problem and applies to settings where the probability measures $\mu$ and $\nu$ are either discrete or continuous [12]. In general, establishing the uniqueness, and structure of the optimal transport map $T$ that solves (2) is challenging [14]. In fact, a map satisfying $T_{\#}\mu=\nu$ may not exist at all since the Monge problem does not permit mass splitting. For example, consider discrete measures where $\mu$ assigns all its mass to a single point, while $\nu$ distributes the mass evenly between two distinct points. Since $T$ is a deterministic map sending each point $x\in\mathcal{X}$ to a single point $y\in\mathcal{Y}$, no such $T$ can push $\mu$ forward to $\nu$ [12].
The existence and uniqueness of the solution to the Monge problem in (2) can be established with additional assumptions on the probability measures and cost function [14]. In this work, we specify $\mathcal{X}=\mathcal{Y}=\mathbb{R}^{N}$ and assume $T$ is a diffeomorphism. [^1] See Theorem 3.6 in [14] for generalizations of this setting. Brenier’s theorem [11, 15] guarantees that if the probability measures $\mu$ and $\nu$ are absolutely continuous with respect to the Lebesgue measure, i.e., they admit respective probability densities $p(x)$ and $q(y)$, and the cost function is squared Euclidean distance, the solution to (7) is unique and can be expressed as
@@
[…]
where $J_{T}(x)$ denotes the $N\times N$ Jacobian of $T$ evaluated at $x$ and where the second equality follows from the change of variables formula applied to the substitution $y=T(x)$. As described in [11, 8], this yields the following standard pushforward expression:
$p(x)=q(T(x))|\det J_{T}(x)|$ (5)
@@
$p(x)$ $=q(\nabla\phi(x))\det(H_{\phi}(x))$ (6)
where $H_{\phi}=J_{T}$ is the $N\times N$ positive semidefinite Hessian of $\phi$.
@@ IV Related Work
Other works have considered parameterizing the monotone gradient map without first characterizing the underlying potential. For example, [22] parameterize a Gram factor of $H_{\phi}$, the positive semidefinite Hessian of the potential $\phi$, and obtain the gradient map by numerically integrating the learned Hessian. The only known architecture suitable for parameterizing the Gram factor is a single-layer network of the form $\sigma(Wx+b)$, which does not require numerical integration and serves primarily as a proof of concept. Moreover, the reliance on numerical integration poses significant challenges for optimal transport in high-dimensional settings.
In contrast, we first proposed neural network architectures that are guaranteed to directly model monotone gradients of convex functions in [1]. Then, in [2], we first proved that these neural networks are universal approximators of monotone gradient functions and then introduced mGradNets, which offer expanded parameterizations more amenable for optimization in practice. In this work, we leverage the Monge-Ampère equation to design a principled approach for training mGradNets to solve Monge optimal transport problems where both the source and target probability densities are known. This contrasts with our approach in [1], where the source probability density was unknown and we trained MGNs (predecessors to mGradNets) by maximizing the target likelihood of the mapped samples – the approach taken previously by [18].
@@ V Proposed Method
[…]
As a result of Brenier’s theorem, the optimal transport map $T$ solving (7) has the following properties: 1) $T$ has a symmetric positive semidefinite Jacobian since it is the monotone gradient of a convex function; 2) $T$ satisfies the pushforward expression in (5) that is closely related to the Monge-Ampère equation. Hence, we propose to parameterize $T$ as a monotone gradient network (mGradNet) [1, 2], and train it to satisfy the constraint in (5) since any function that satisfies the constraint and is the gradient of a convex function is guaranteed to be the optimal transport map.
@@
Figure 1: OT from skewed to standard Gaussian.
mGradNets are a class of neural network architectures that are explicitly designed to correspond to monotone gradients of convex functions. The Jacobian of a mGradNet’s output with respect to its input is always symmetric positive semidefinite – for all inputs $x$ and all parameter values $\theta$ – thereby guaranteeing that for any mGradNet $T_{\theta}$, there exists a convex potential $\phi_{\theta}$ such that $T_{\theta}=\nabla\phi_{\theta}$. Moreover, mGradNets are universal approximators of gradients of convex functions and can be additionally tailored to correspond to subclasses of convex functions, such as convex functions expressible as the sum of convex ridge functions. We refer to our paper [2] for further details and proofs regarding mGradNets and GradNets, their generalization to gradients of $L$-smooth functions.
@@
mGradNet $T_{\theta}$, source density $p(x)$, target density $q(y)$, learning rate $\eta$, batch size $B$, iterations $I$
for iteration $i=1$ to $I$ do
     Sample $\{x_{1},x_{2},\dots,x_{B}\}\sim p(x)$
     Compute pushforward points: $y_{j}=T_{\theta}(x_{j})$
     Compute Jacobians: $J(\theta,x_{j})=\frac{\partial}{\partial x_{j}}T_{\theta}(x_{j})$
@@
$\mathcal{L}(\theta)=\frac{1}{B}\sum_{j=1}^{B}\ell\left(\log\det J(\theta,x_{j}),\log p(x_{j})-\log q(y_{j})\right)$
     Update parameters: $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}(\theta)$
@@
return $\theta$
The training procedure, as provided in Algorithm 1, is inspired by the training of physics informed neural networks [24] designed to satisfy differential equations. During training, points $x_{j}$ are sampled from the source distribution $p$ and transformed by the mGradNet $T_{\theta}$. Evaluating the density of the resulting transformed points using the target density $q$ leads to computing the training label as $\log p(x_{j})-\log(q(T_{\theta}(x_{j}))$. The log determinant of the Jacobian of $T_{\theta}$ (with respect to the input) is evaluated at $x_{j}$ and trained to match this label. The training procedure presented in Algorithm 1 minimizes the empirical version of the following objective over the mGradNet parameters $\theta$, where $\ell$ is a loss function like the squared Euclidean distance function:
$\mathbb{E}_{x\sim p(x)}\ell(J_{T_{\theta}}(x),\log p(x)-\log(q(T_{\theta}(x)))$ (8)
@@ VI Experiments
We demonstrate the efficacy of the proposed GradNetOT procedure with a series of optimal transport tasks. In all experiments, we use Adam optimizer and a fixed batch size of 1000.
@@ VI-A Gaussians
We consider the OT problem in (7) with quadratic cost and Gaussian measures $\mu$ and $\nu$. The OT map between multivariate Gaussians $\mathcal{N}(\mu_{1},\Sigma_{1})$ and $\mathcal{N}(\mu_{2},\Sigma_{2})$ has an analytic solution. In particular, when the target density is the multivariate standard normal, the OT map is given by the whitening transform $T(x)=\Sigma_{0}^{-1/2}(x-\mu_{0})$. We train the mGradNet-C and mGradNet-M architectures from [2] using Algorithm 1. Using the same method, we also a train a baseline feedforward neural network that is not constrained to be the gradient of a convex function. We visualize the OT maps for the neural network baseline, mGradNet-C, mGradNet-M and the ground truth OT map whitening in Fig.1. The points are colored in the figure according to their corresponding positions in the source distribution. As shown in the figure, the neural network baseline fails to learn the OT map as indicated by the fact that the ordering of the points changes drastically: the red source points are on the right in Fig.1(a), but they are mapped to the red points on the left in Fig.1(b). This indicates that the baseline neural network is not learning a monotone mapping. Meanwhile, the mGradNet-C and mGradNet-M can learn the OT map and the displacements closely match those of the known OT map, namely the whitening transform.
To test with higher-dimensional Gaussians, we randomly generate the $d$-dimensional source Gaussian distribution by sampling a mean from the standard Gaussian distribution and by sampling the covariance matrix from a Wishart distribution with $d+1$ degrees of freedom and identity scale matrix. After training mGradNets using Algorithm 1, we evaluate them on a set of 1000 test points. The whitening transform applied to the test points yields the ground truth target under the optimal transport map. We use the mGradNets to map the test points and compute the mean squared between them and the ground truth mapped points. For all dimensions, both mGradNets closely approximate the optimal transport map and the mGradNet-M consistently outperforms the mGradNet-C.
@@
VI-B Robot Swarms as Optimal Transport
We consider the problem of controlling a swarm of UAVs or robots from their known initial locations to feasible desired locations. We model the swarm arrangements as probability densities and, without loss of generality, set the initial time to $t=0$ and the target time to $t=1$. We consider deployment settings where a user is concerned with the swarm’s overall arrangement rather than specific locations of individual robot – such as a drone light show where a user inputs both initial and desired swarm configurations using grayscale images.
This leads to an optimal transport problem for arbitrary, non-Gaussian densities in $\mathbb{R}^{2}$. Specifically, for this experiment, we model image pixel intensities from the MNIST dataset as samples from a continuous density over the unit square $[0,1]^{2}$. Each pixel is treated as a point mass located on a uniform grid, and we fit a smooth density using kernel density estimation (KDE). Given an $n\times n$ image with intensity values $I(i,j)$, we the density over pixel locations $x$ is:
$p(x)=\sum_{i=0}^{n-1}\sum_{j=0}^{n-1}I(i,j)\cdot\mathcal{N}\left(x;\left(\frac{i}{n-1},\frac{j}{n-1}\right),\sigma^{2}I\right)$ (9)
with $\sigma=1\times 10^{-4}$ and $n=28$ for MNIST. We train an mGradNet-M to learn the optimal transport map between these densities using Algorithm 1. During training, we exponentially decay the learning rate from 1e-2 to 1e-4 over 10000 iterations and also decay $\sigma$ in the same manner to enable more stable training. The results for morphing between a 0 and a 4, using the displacement map $x=(1-t)x+tT_{\theta}(x)$ for $t=\{0,0.25,0.5,0.75,1\}$, which gives the geodesic between distributions according to the Wasserstein-2 metric [25], are shown in Fig. 3.

### v2

@@ Abstract
Monotone gradient functions play a central role in solving the Monge formulation of the optimal transport (OT) problem, which arises in modern applications ranging from fluid dynamics to robot swarm control. When the transport cost is the squared Euclidean distance, Brenier’s theorem guarantees that the unique optimal transport map satisfies a Monge-Ampère equation and is the gradient of a convex function. In [1, 2], we proposed Monotone Gradient Networks (mGradNets), neural networks that directly parameterize the space of monotone gradient maps. In this work, we leverage mGradNets to directly learn the optimal transport mapping by minimizing a training loss function defined using the Monge-Ampère equation. We empirically show that the structural bias of mGradNets facilitates the learning of optimal transport maps across both image morphing tasks and high-dimensional OT problems.
@@ I Introduction
The optimal transport (OT) problem originated with the work of Gaspard Monge in 1781, who was interested in a fundamental problem that arises when constructing fortifications: how can one transfer molecules of soil from piles to holes while minimizing the cost of transfer [3]? Since then, Monge’s optimal transport problem and Leonid Kantorovich’s relaxation [4, 5] have been phrased in terms of probability densities and finding an optimal transport map: how can a source probability distribution be mapped to a target probability distribution in a way that minimizes the transport cost?
These formulations of optimal transport, along with several newer variations, have found applications spanning not only statistics and logistics, but also signal and image processing, robotics, machine learning, fluid dynamics, economics, geology, astronomy, and beyond. In our simulations, we illustrate an example of optimal transport for image morphing applications. Robotics applications of OT include the control of large swarms of robots or unmanned aerial vehicles (UAVs) [6], which have applications ranging from critical natural disaster rescue missions to intricate and artistic light shows. By modeling movements of large robot swarms as time-evolving spatial probability distributions, an optimal transport map can guide the swarm from an initial distribution to a desired target distribution while minimizing travel cost [7]. Furthermore, OT applications to economics include works showing the equivalence between optimal transport, stable matching, and modeling hedonic price equilibria, which capture relationships between prices and product characteristics [8]. In geology, developments concerning OT on Riemannian manifolds have been used to model continental drift by describing ground mass distributions and geodesic transport paths over time [9]. In astrophysics, recent works on Monge-Ampère gravity employ optimal transport maps to model the distribution of mass throughout the universe and the formation of cosmological structures [10].
@@
Exploiting Brenier’s link between optimal transport and the Monge-Ampère equation is nontrivial because the Monge-Ampère equation is highly non-linear and generally difficult to solve using standard techniques like calculus of variations [12]. While some works, such as [10], rely on additional assumptions like spherical symmetry to recover an analytic solution, the majority of the literature uses numerical approaches like discrete-discrete entropic regularization and semi-discrete optimal transport [13]. In this paper, we focus on the intersection of Brenier’s result and advancements in neural networks to solve Monge optimal transport problems.
We propose GradNetOT: a method that leverages monotone gradient networks (mGradNets) [2] to solve Monge optimal transport problems. Using an mGradNet to parameterize the optimal transport map, we search through the space of all gradients of convex potential functions. To guide our search, we explicitly use the Monge-Ampère equation in our optimization objective.
@@ II Background
We briefly review probability concepts relevant to optimal transport (OT). A probability measure $\mu$ assigns mass to events, which are measurable subsets of a sample space $\mathcal{X}$. The probability measure satisfies the Kolmogorov axioms of non-negativity, countable additivity, and normalization ($\mu(\mathcal{X})=1$). The probability of an event $A\subseteq\mathcal{X}$ is expressed as $\mu(A)=\int_{A}d\mu({\mathbf{x}})$. For example, if $\mu$ is discrete, it assigns probability mass to a finite set of points and $\mu(A)=\sum_{x\in A}\mu({\mathbf{x}})$. Alternatively, if $\mathcal{X}=\mathbb{R}^{N}$, then the Lebesgue measure yields the $n$-dimensional volume of $A\subset\mathbb{R}^{N}$. If $\mu$ is absolutely continuous with respect to the Lebesgue measure, then it admits a probability density function $p({\mathbf{x}})$ and $\mu(A)=\int_{A}p({\mathbf{x}})dx$.
@@
$f_{\#}\mu(A)=\mu(f^{-1}(A))=\mu(\{{\mathbf{x}}\in\mathcal{X}:f({\mathbf{x}})\in A\})$ (1)
where $f^{-1}(A)$ corresponds to the preimage of $A$ under $f$. Intuitively, (1) computes the probability of set $A$ in the target space $\mathcal{Y}$ by measuring the probability of the set in the source space $\mathcal{X}$ whose image under $f$ is $A$. Additionally, $f$ is called a diffeomorphism if $f$ is a bijection and both $f$ and $f^{-1}$ are differentiable.
@@ III Problem Statement
Optimal transport involves finding the most efficient way to transform one probability distribution into another. The distributions can be thought of as descriptions of how mass (representing probability) is spread over a space. Efficiency here is measured by a predefined cost function.
Formally, let $\mathcal{X}$ and $\mathcal{Y}$ be spaces equipped with respective probability measures $\mu$ and $\nu$. Let $c:\mathcal{X}\times\mathcal{Y}\to[0,\infty)$ be a cost function, where $c({\mathbf{x}},{\mathbf{y}})$ denotes the cost of moving a unit of mass (probability) from $x\in\mathcal{X}$ to $y\in\mathcal{Y}$. The optimal transport problem is then to find a map $T:\mathcal{X}\to\mathcal{Y}$ that pushes forward $\mu$ to $\nu$ (i.e., $T_{\#}\mu=\nu$), and minimizes the total transport cost:
$\inf_{T:T_{\#}\mu=\nu}$ $\int_{\mathcal{X}}c({\mathbf{x}},T({\mathbf{x}}))d\mu({\mathbf{x}})$ (2)
The formulation in (2) is known as the Monge problem and applies to settings where the probability measures $\mu$ and $\nu$ are either discrete or continuous [12]. In general, establishing the uniqueness, and structure of the optimal transport map $T$ that solves (2) is challenging [14]. In fact, a map satisfying $T_{\#}\mu=\nu$ may not exist at all since the Monge problem does not permit mass splitting. For example, consider discrete measures where $\mu$ assigns all its mass to a single point, while $\nu$ distributes the mass evenly between two distinct points. Since $T$ is a deterministic map sending each point ${\mathbf{x}}\in\mathcal{X}$ to a single point ${\mathbf{y}}\in\mathcal{Y}$, no such $T$ can push $\mu$ forward to $\nu$ [12].
The existence and uniqueness of the solution to the Monge problem in (2) can be established with additional assumptions on the probability measures and cost function [14]. In this work, we specify $\mathcal{X}=\mathcal{Y}=\mathbb{R}^{N}$ and assume $T$ is a diffeomorphism. [^1] See Theorem 3.6 in [14] for generalizations of this setting. Brenier’s theorem [11, 15] guarantees that if the probability measures $\mu$ and $\nu$ are absolutely continuous with respect to the Lebesgue measure, i.e., they admit respective probability densities $p({\mathbf{x}})$ and $q({\mathbf{y}})$, and the cost function is squared Euclidean distance, the solution to (7) is unique and can be expressed as
@@
$\int_{T^{-1}(A)}p({\mathbf{x}})d{\mathbf{x}}$ $=\int_{A}q({\mathbf{y}})d{\mathbf{y}}=\int_{T^{-1}(A)}q(T({\mathbf{x}}))|\det{\mathbf{J}}_{T}({\mathbf{x}})|d{\mathbf{x}}$ (4)
where ${\mathbf{J}}_{T}(x)$ denotes the $N\times N$ Jacobian of $T$ evaluated at $\mathbf{x}$} and where the second equality follows from the change of variables formula applied to the substitution ${\mathbf{y}}=T({\mathbf{x}})$. As described in [11, 10], this yields the following standard pushforward expression:
$p({\mathbf{x}})=q(T({\mathbf{x}}))|\det{\mathbf{J}}_{T}({\mathbf{x}})|$ (5)
@@
$p({\mathbf{x}})$ $=q(\nabla\phi({\mathbf{x}}))\det({\mathbf{H}}_{\phi}({\mathbf{x}}))$ (6)
where ${\mathbf{H}}_{\phi}={\mathbf{J}}_{T}$ is the $N\times N$ positive semidefinite Hessian of $\phi$.
@@ IV Related Work
Other works have considered parameterizing the monotone gradient map without first characterizing the underlying potential. For example, [22] parameterize a Gram factor of ${\mathbf{H}}_{\phi}$, the positive semidefinite Hessian of the potential $\phi$, and obtain the gradient map by numerically integrating the learned Hessian. The only known architecture suitable for parameterizing the Gram factor is a single-layer network of the form $\sigma({\mathbf{W}}{\mathbf{x}}+{\mathbf{b}})$, which does not require numerical integration and serves primarily as a proof of concept. Moreover, the reliance on numerical integration poses significant challenges for optimal transport in high-dimensional settings.
In contrast, we proposed neural network architectures that are guaranteed to directly model monotone gradients of convex functions in [1]. Then, in [2], we proved that these neural networks are universal approximators of monotone gradient functions and introduced mGradNets that offer expanded parameterizations more amenable for optimization in practice. In this work, we leverage the Monge-Ampère equation to design a principled approach for training mGradNets to solve Monge optimal transport problems where both the source and target probability densities are known. This contrasts with our approach in [1], where the source probability density was unknown and we trained MGNs (predecessors of mGradNets) by maximizing the target likelihood of the mapped samples – the approach taken previously by [18].
@@ V Proposed Method
$\inf_{T:T_{\#}\mu=\nu}$ $\int_{\mathcal{X}}\|{\mathbf{x}}-T({\mathbf{x}})\|^{2}d\mu({\mathbf{x}})$ (7)
As a result of Brenier’s theorem, the optimal transport map $T$ solving (7) has the following properties: 1) $T$ has a symmetric positive semidefinite Jacobian since it is the monotone gradient of a convex function; 2) $T$ satisfies the pushforward expression in (5) that is closely related to the Monge-Ampère equation. Hence, we propose to parameterize $T$ as a monotone gradient network (mGradNet) [1, 2], and train it to satisfy the constraint in (5). This leverages Brenier’s theorem [11]: a function that is the gradient of a convex function and satisfies the constraint must be the unique optimal transport map.
@@
Figure 1: OT from skewed to standard Gaussian. Points are colored based on their positions in the source distribution.
mGradNets are a class of neural network architectures that are explicitly designed to correspond to monotone gradients of convex functions. The Jacobian of a mGradNet’s output with respect to its input is always symmetric positive semidefinite – for all inputs $\mathbf{x}$} and all parameter values $\boldsymbol{\theta}$ – thereby guaranteeing that for any mGradNet $T_{\boldsymbol{\theta}}$, there exists a convex potential $\phi_{\boldsymbol{\theta}}$ such that $T_{\boldsymbol{\theta}}=\nabla\phi_{\boldsymbol{\theta}}$. Moreover, mGradNets are universal approximators of gradients of convex functions and can be additionally tailored to correspond to subclasses of convex functions, such as convex functions expressible as the sum of convex ridge functions. We refer to our paper [2] for further details and proofs regarding mGradNets and GradNets (a generalization to gradients of $L$-smooth functions).
@@
mGradNet $T_{\boldsymbol{\theta}}$, source density $p({\mathbf{x}})$, target density $q({\mathbf{y}})$, learning rate $\eta$, batch size $B$, iterations II
for iteration $i=1$ to II do
  Sample $\{{\mathbf{x}}_{1},{\mathbf{x}}_{2},\dots,{\mathbf{x}}_{B}\}\sim p({\mathbf{x}})$
  Compute pushforward points: ${\mathbf{y}}_{j}=T_{\boldsymbol{\theta}}({\mathbf{x}}_{j}),\;j=\{1,\dots,B\}$
  Compute Jacobians: ${\mathbf{J}}_{T_{\boldsymbol{\theta}}}({\mathbf{x}}_{j})=\frac{\partial}{\partial{\mathbf{x}}_{j}}T_{\boldsymbol{\theta}}({\mathbf{x}}_{j})$
@@
$\mathcal{L}(\boldsymbol{\theta})=\frac{1}{B}\sum_{j=1}^{B}\ell\left(\log\det{\mathbf{J}}_{T_{\boldsymbol{\theta}}}({\mathbf{x}}_{j}),\log p({\mathbf{x}}_{j})-\log q({\mathbf{y}}_{j})\right)$
  Update parameters: $\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}-\eta\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta})$
@@
return $\boldsymbol{\theta}$
The training procedure, as provided in Algorithm 1, is inspired by the training of physics informed neural networks [24] designed to satisfy differential equations. During training, points $\{{\mathbf{x}}_{j}\}_{j=1}^{B}$ are sampled from the source distribution $p$ and transformed by the mGradNet $T_{\boldsymbol{\theta}}$. Using the target density $q$ to evaluate the density of the transformed points leads to computing the training label as $\log p({\mathbf{x}}_{j})-\log(q(T_{\boldsymbol{\theta}}({\mathbf{x}}_{j}))$. The log determinant of the Jacobian of $T_{\boldsymbol{\theta}}$ (with respect to the input) is evaluated at ${\mathbf{x}}_{j}$ and trained to match this label via standard backpropagation. The training procedure presented in Algorithm 1 minimizes the empirical version of the following objective over the mGradNet parameters $\boldsymbol{\theta}$, where $\ell$ is a loss function like the squared Euclidean distance function:
$\mathbb{E}_{{\mathbf{x}}\sim p({\mathbf{x}})}\ell({\mathbf{J}}_{T_{\boldsymbol{\theta}}}({\mathbf{x}}),\log p({\mathbf{x}})-\log(q(T_{\boldsymbol{\theta}}({\mathbf{x}})))$ (8)
@@ VI Experiments
We demonstrate the efficacy of the proposed GradNetOT procedure with a series of optimal transport tasks. In all experiments, we use the Adam optimizer and a fixed batch size of 1000.
@@ VI-A Gaussians
We consider the OT problem in (7) with quadratic cost and Gaussian measures $\mu$ and $\nu$. The OT map between multivariate Gaussians $\mathcal{N}(\boldsymbol{\mu}_{0},\boldsymbol{\Sigma}_{0})$ and $\mathcal{N}(\boldsymbol{\mu}_{1},\boldsymbol{\Sigma}_{1})$ has an analytic solution. In particular, when the target density is the multivariate standard normal ($\boldsymbol{\mu}_{1}=\boldsymbol{0},\boldsymbol{\Sigma}_{1}=\boldsymbol{I}$), the OT map is given by the whitening transform $T({\mathbf{x}})=\boldsymbol{\Sigma}_{0}^{-1/2}({\mathbf{x}}-\boldsymbol{\mu}_{0})$. We train the mGradNet-C and mGradNet-M architectures from [2] using Algorithm 1. Using the same method, we also a train a baseline feedforward neural network that is not constrained to be the gradient of a convex function. We visualize the OT maps for the neural network baseline, mGradNet-C, mGradNet-M and the ground truth OT map whitening in Fig. 1. The points are colored in the figure according to their corresponding positions in the source distribution. As shown in the figure, the neural network baseline fails to learn the OT map as indicated by the fact that the ordering of the points changes drastically: the red source points are on the right in Fig. 1(a), but they are mapped to the red points on the left in Fig. 1(b). This indicates that the baseline neural network is not learning a monotone mapping. Meanwhile, the mGradNet-C and mGradNet-M can learn the OT map and their displacements closely match those of the known OT map, namely the whitening transform.
To test with higher-dimensional Gaussians, we randomly generate a $d$-dimensional source Gaussian distribution by sampling the mean from the standard Gaussian distribution and by sampling the covariance matrix from a Wishart distribution with $d+1$ degrees of freedom and identity scale matrix. After using Algorithm 1 to train mGradNets to learn the OT map to a standard Gaussian, we evaluate the neural networks on a set of 1000 test points. The whitening transform applied to the test points again yields the ground truth target under the optimal transport map. We use our mGradNets to map the test points and compute the mean squared between them and the ground truth mapped points. Fig. 2 shows that, for all dimensions, both mGradNets closely approximate the optimal transport map and the mGradNet-M consistently outperforms the mGradNet-C.
@@
VI-B Image Morphing as Optimal Transport
We consider the problem of optimally transporting a distribution over pixel locations from their known initial locations to desired locations, thereby effectively morphing the image. We model image pixel intensities from the MNIST dataset as samples from a continuous density over the unit square $0,1]^{2}$. Each pixel is treated as a point mass located on a uniform grid and we fit a smooth density using kernel density estimation (KDE). Given an $n\times n$ image with intensity values $I(i,j)$, we model the density over pixel locations ${\mathbf{x}}\in[0,1]^{2}$ as the following mixture of Gaussians:
$p({\mathbf{x}})=\sum_{i=0}^{n-1}\sum_{j=0}^{n-1}I(i,j)\cdot\mathcal{N}\left({\mathbf{x}};\left(\frac{i}{n-1},\frac{j}{n-1}\right),\sigma^{2}\mathbf{I}\right)$ (9)
with $\sigma^{2}=10^{-4}$ and $n=28$ for MNIST. We train an mGradNet-M to learn the optimal transport map between these densities using Algorithm 1. During training, we exponentially decay the learning rate from 1e-2 to 1e-4 over 10000 iterations and also decay $\sigma^{2}$ in the same manner to facilitate more stable training. In Fig. 3, we illustrate the learned OT displacement map
${\mathbf{x}}^{(t)}=(1-t){\mathbf{x}}^{(0)}+tT_{\boldsymbol{\theta}}\left({\mathbf{x}}^{(0)}\right)$ (10)
for morphing between a 0 and a 4.